# BroadGen: A Framework for Generating Effective and Efficient Advertiser Broad Match Keyphrase Recommendations

### v3

@@ I Introduction
Advertisers or sellers on e-commerce platforms are recommended keyphrases by the platform to increase their item visibility depending on buyer searches as shown in Figure 1(a). The sellers bid on relevant recommended keyphrases as well as their own custom keyphrases to increase their sales (Figure 1(b)). Keyphrase recommendation is the subject of a lot of research and well studied in several works [1, 2, 3, 4, 5, 6, 7], due to their challenging research and industrial importance. We term the recommendation to the seller as Keyphrase and the buyer search text as Query. Additionally, the platform matches the (recommended) keyphrases bid by the sellers to the buyer search queries [8].
(a) Buyer side
(b) Seller Side
Figure 1: Screenshot of keyphrases for manual targeting in Promoted Listings Priority for eBay Advertising.
@@
Previous research has focused on recommending exact match type keyphrases, popularly framing it as an Extreme Multi-Label Classification (XMC) [9, 10], where given an input item, they determine relevance to the labels which in this case are the buyer queries or keyphrases. Match types such as phrase and broad match keyphrases can alleviate problems associated with exact match type recommendations (see Section I-A). However, their recommendation strategies are under-researched as they are unlike exact match recommendations. Therefore, in this paper, we focus on other match type keyphrases by formulating the problem, designing solutions and providing insights learned in production.
Figure 2: Matching architecture
@@ I-A1 Shifting label distributions
Analysis reveals that in sponsored search, queries adhere to the 80/20 rule where 20% of queries generate 80% of search volume leaving 80% as long-tail queries [11], which is prominent even in e-commerce platforms like eBay. These long-tail queries are low in volume, transient, and rarely reappearing monthly. This transient nature limits the effectiveness of traditional exact match algorithms [1, 2, 3, 4, 12], as most labels won’t have future search volume. Due to the shifting dynamics of long-tail queries, predicting them individually as exact match keywords is nearly infeasible at time of inference. Instead, using other match types addresses this issue without explicit demand forecasting. Broad match keyphrases inherently address this challenge by capturing a wider set of related queries, thus effectively mitigating the unpredictability and transient nature of long-tail search patterns.
@@ I-A2 Management costs and Sales
Managing sponsored search campaigns demands substantial effort from sellers, compounded by the extensive set of keyphrases requiring bids [11, 13, 14]. Sellers prefer concise recommendations due to their perception and business demands [2, 12]. Other match type keyphrases are beneficial for targeting a wide array of exact queries, thereby reducing seller costs and involvement. Research suggests that when curation accuracy is high and seller bids remain competitive, broad match keyphrases yield higher sales [14].
@@ I-B1 Ground Truth Hurdles
Exact match keyphrases can be directly recommended by modeling on click data logs, since they correspond precisely to buyer queries. In contrast, phrase and broad match keyphrases are more difficult to model due to the sparsity and bias of available training signals. It suffers from sparsity as the performance information of phrase/broad match types are only collected when the seller manually chooses them. Although the sellers’ previous picks of successful broad match keyphrases could seem like a good basis for ground truth, they can be heavily affected by the seller set bid amounts and auctioning dynamics [15], thus making seller-based preference data a bad choice for modeling keyphrase recommendations [12].
Several biases with click data have been well studied [2, 16, 17, 18, 8], these impact any form of supervised training for recommendation. The limited availability of ground truth data leaves a significant portion of the feature space unmarked, requiring the incorporation of counterfactual data so models can learn significant patterns. Moreover, missing label biases [19] affects substantial counterfactual negative signals, as the absence of impressions or clicks doesn’t mean that the match type keyphrases are irrelevant.
@@ I-B2 Active Targeting
Match type keyphrases recommended to sellers should be reachable to queries that are searched by buyers. Generated broad/phrase match keyphrases which are open vocabulary cannot reach such queries, though relevant to the item are considered futile leading to wastage of seller bids. [^1] See discussion of Out-of-Vocabulary models [20, 21, 22, 5] and targeting in Graphex [2]. This is exacerbated by the transient queries mentioned in Section I-A1 that require keyphrases to target queries that buyers search for over time. Thus, there should a way of evaluating performance of broad matches on transient queries.
@@ I-B4 Execution Requirements
Executing billions of item and query inferences within a few hours is paramount for the timely completion of daily tasks. Given the high memory demands, single-node data aggregation for inference poses a failure risk, while multi-node management elevates engineering expenses. Models need frequent updates for continuous integration of new data points, as discussed in Section I-B2. Deep learning networks like Large Language Models (LLMs) face scrutiny for high execution costs, notably due to their training and inference latency [23, 2, 24, 25]. Leading LLMs often rely on GPUs, raising acquisition costs and affecting both advertisers’ and platforms’ profit margins.
@@ I-C Scope and Contributions
In this work we only focus on Broad matches instead of Phrase matches  [26, 27] due to their wide applicability. We use our definition in Section I to streamline the evaluation; the actual implementation of the matching is platform-specific. We aim to tackle the challenges mentioned in Section I-B, which we briefly describe here. We present a novel framework BroadGen for recommending broad match keyphrases to sellers and evaluate its performance. Due to the lack of ground truths per item (see Challenge I-B1), we incorporate relevant buyer queries as input (see Section IV-A) to determine the effective keyphrases for an item. This also helps us target active queries for Challenge I-B2. The Challenge I-B3 implies that a good broad match keyphrase should be neither too generic nor too specific, so that it can achieve a balance between reach and effectiveness. A “too specific” broad match keyphrase will behave like an exact match keyphrase which defeats the purpose. Our approach emphasizes avoiding overly broad suggestions, particularly in Section III-E, with the aim of mitigating inefficiencies within both our systems and the business metrics of sellers. Our framework has been designed to be light and fast with minimal infrastructure costs and is easily deployable in a distributed manner (Challenge I-B4) to cater to millions of sellers with an inventory of over 2.5 billion items daily. We also develop a unique evaluation methodology described in Sections IV-A and  IV-B that enables us to calculate the balance in effectiveness and efficiency of the recommended broad match keyphrases in relation to the search queries it will be matched to. It allows for absence of ground truths and assesses performance over time, emphasizing their adaptability and temporal resilience (Challenge I-B2).
@@ II Related Work
Although broad match keyphrases are a crucial component, the research dedicated to specifically recommending novel broad match keyphrases is sparse. Traditionally, sellers have been responsible for selecting match types, with prior studies [28] concentrating on creating keyphrases suitable for exact matches. A simplistic recommendation strategy has involved a rule-based Match Type Curation, which applies heuristics to assess whether an exact match type keyphrase, obtained through the current recall mechanism, has historically succeeded as a broad match keyphrase. This approach, however, faces intrinsic challenges: without historical data indicating that a seller has previously chosen the keyphrase as a broad match type — this method is unable to suggest the keyphrase due to a lack of pertinent performance indicators. Note that in this method the keyphrase is a query in itself.
Figure 3: Transfusion Algorithm Explained
@@
Some studies have focused on optimizing the matching of key-phrases to queries rather than effecting the sellers/advertiser’s selection of keyphrases. The BB-KSM model [29] is a stochastic model that utilizes Markov Chain Monte Carlo (MCMC) to initially estimate keyphrase metrics such as impressions and CTR across match types, filling in incomplete data. Subsequently, it designates a match type to each keyphrase based on these metrics with the aim of maximizing a profit objective under budgetary constraints, employing a branch-and-bound algorithm for efficient decision space exploration. Similar optimization strategies have been documented for match types [30], which recommend using buyer search queries as broad match keyphrases rather than introducing new broad keyphrases. However, with the increasing variety of items and queries in e-commerce, this approach may be inefficient and expensive. The introduction of new queries and changing distributions (refer to Section I-A1) further reduces profitability. Studies like [31] apply query substitution methods [32, 33] to connect broad keyphrases with buyer queries by preprocessing frequently searched queries and utilizing predefined substitution mappings.
@@
Transfusion is an eBay’s match type generative model developed based on its in-house NER (Named Entity Recognition) model [34, 35] to recommend broad match type keyphrases especially for cold start items. The model generates keyphrases for a specified item utilizing the tokens present in its title. It ensures that these keyphrases align with popular and historically high-performing NER tag combinations within a particular meta category. For instance, as shown in Figure 3, the model identifies that the NER tag combination of product_name and storage_capacity is popular in this category based on the scoring function $f(X)$. The algorithm then construct 3 keyphrases: ‘‘iPhone 5S 16GB’’, ‘‘iPhone 5S 32GB’’, ‘‘iPhone 5S 64GB’’ based on the tokens corresponding to this NER tag combination from the item title. In this work, for comparison, we use a proxy for the rule-based match type curation and the Transfusion model against our model BroadGen.
@@ III-A Terms, Notations and Formulation
Our framework requires a set of existing (buyer search) queries denoted as $Q$ and the item title $t$ for each item to recommend a set of keyphrases denoted as $K$. The goal is for the keyphrase set $K$ to be broad matched to the queries $Q$. Each query $q\in Q$, each keyphrase $k\in K$ and the title $t$ are in the form a list of tokens. [^2] A proprietary tokenization scheme handles normalization and stop word removal. We formulate the problem as String Clustering on the set of query strings ($Q$), to exploit the commonality within a cluster of queries and generate a representation for each cluster. The underlying intuition is that the cluster representation serves as the broad match keyphrase capable of addressing both the queries within the set and prospective queries beyond it. Our framework has three cores, Input generation (III-B) which generates the input matrix from the set of input queries. The Clustering core (III-C) which groups the queries using the input matrix. The Representation generation (III-D) constructs a representative ordered string for each cluster.
@@ III-B Input Generation
Two input matrices are constructed, Anchor Similarity Matrix $A^{|Q|\times|t|}$ and Query-Query Similarity Matrix $S^{|Q|\times|Q|}$, where $|\cdot|$ determines cardinality. In constructing $A$, a $|t|$-dimensional vector is generated for each query. This vector comprises of binary values, where 1s and 0s indicate the query token’s presence in the title. The construction of $S$ is done by computing the common token count between each pair of queries in $Q$. Each row of $S$ can be considered as a $|Q|$-dimensional vector that represents each query depending on similarity with other queries. Each vector of matrix $S$ is normalized by dividing by its maximum, while each vector of the matrix $A$ is normalized by its sum. The Query-Query Similarity matrix $S$ is transformed into a distance matrix by $D_{ij}=1-S_{ij}$.
@@
The core combines both the matrices $A$ and $S$ as shown in Equation 1. The normalized augmented matrix $\hat{Z}$, provides a vector representation for each query anchoring in relation to the title and acts as a query distance vector when there is insufficient information in the anchor representation.
@@ III-C Clustering Core
The core functionality starts with computing the pairwise distance between each pair of vectors $z_{i},z_{j}\in\hat{Z}$ as demonstrated in Equation 2. Now, with the pairwise distance matrix $\tilde{Z}$, the core performs agglomerative clustering [36] by constructing linkage distances with ward variance minimization [37] and optimal leaf ordering [38]. We denote the clustering function as $R(\cdot,\cdot)$.
@@ III-D Representation Generation
A representation for each cluster in $C$ depicts a keyphrase that can be extended to queries in the cluster. In the final core, the representation is determined as the set of tokens common among the tokenized queries within each cluster. At times, clusters can be large leading to single or no tokens (in absence of commonality). In these situations, an optimal subset of the cluster is identified whose representation is more than one token. The common tokens from each cluster need to be ordered before concatenating into a string. We employ a graph-based approach to infer an ordering of the common tokens based on their relative positions across the queries. A graph is constructed with all tokens as the vertices; for each co-occurring pair of tokens a directed edge is added indicating local precedence. A topological sort [39] of the graph using a depth-first search algorithm produces a consistent global ordering of the tokens. This method of determining precedence is rooted in prior work, such as sentence ordering task [40]. For each cluster, the common tokens are arranged following the global order and represented as a string.
Figure 4: Illustration of BroadGen.
@@ IV-A1 OOT Formulation
To evaluate the effectiveness of broad match keyphrases in addressing the challenge of shifting label distribution, we assess their generalized performance in an Out-of-Time (OOT) [20] setting.  [^5] Commonly referred to as out-of-distribution generalization [41, 42, 43]. Specifically, the algorithms for generating broad match keyphrases use as input the historical queries searched by buyers prior to time $T$, termed as Pre-queries and evaluated against the query distribution observed in the subsequent $T+30$ day period called as Post-queries.

### v5

@@ I Introduction
Advertisers or sellers on e-commerce platforms are recommended keyphrases by the platform to increase their item visibility depending on buyer searches as shown in Figure 1(a). The sellers bid on relevant recommended keyphrases as well as their own custom keyphrases to increase their sales (Figure 1(b)). Keyphrase recommendation is the subject of a lot of research and well studied in several works [unidec_www25, mishra2025graphexgraphbasedextractionmethod, ashirbad-etal-2024, dahiya2021deepxml, triangular_bidgen, renee_2023, Dahiya23b], due to their challenging research and industrial importance. We term the recommendation to the seller as Keyphrase and the buyer search text as Query. Additionally, the platform matches the (recommended) keyphrases bid by the sellers to the buyer search queries [dey2025middlemanbiasadvertisingaligning].
(a) View of Buyer side.
(b) View of Seller side.
Figure 1: A view of buyer (a) and seller side (b) of the keyphrase recommendation system for eBay advertising (manual targeting).
@@
Previous research has focused on recommending exact match type keyphrases, popularly framing it as an Extreme Multi-Label Classification (XMC) [Bhatia16, agrawal2013multi], where given an input item, they determine relevance to the labels which in this case are the buyer queries or keyphrases. Match types such as phrase and broad match keyphrases can alleviate problems associated with exact match type recommendations (see Section I-A). However, their recommendation strategies are under-researched as they are unlike exact match recommendations. Therefore, in this paper, we focus on other match type keyphrases by formulating the problem, designing solutions and providing insights learned in production.
Figure 2: Architecture for matching seller’s keyphrases with different match types to buyer queries.
@@ I-A1 Shifting label distributions
Analysis reveals that in sponsored search, queries adhere to the 80/20 rule where 20% of queries generate 80% of search volume leaving 80% as long-tail queries [ZHANG2014508], which is prominent even in e-commerce platforms like eBay. These long-tail queries are low in volume, transient, and rarely reappearing monthly. This transient nature limits the effectiveness of traditional exact match algorithms [unidec_www25, mishra2025graphexgraphbasedextractionmethod, ashirbad-etal-2024, dahiya2021deepxml, dey2025judgejudgeusingllm], as most labels won’t have future search volume. Due to the shifting dynamics of long-tail queries, predicting them individually as exact match keywords is nearly infeasible at time of inference. Instead, using other match types addresses this issue without explicit demand forecasting. Broad match keyphrases inherently address this challenge by capturing a wider set of related queries, thus effectively mitigating the unpredictability and transient nature of long-tail search patterns.
@@ I-A2 Management costs and Sales
Managing sponsored search campaigns demands substantial effort from sellers, compounded by the extensive set of keyphrases requiring bids [ZHANG2014508, Amaldoss-2016, YANG2021113491]. Sellers prefer concise recommendations due to their perception and business demands [mishra2025graphexgraphbasedextractionmethod, dey2025judgejudgeusingllm]. Other match type keyphrases are beneficial for targeting a wide array of exact queries, thereby reducing seller costs and involvement. Research suggests that when curation accuracy is high and seller bids remain competitive, broad match keyphrases yield higher sales [YANG2021113491].
@@ I-B1 Ground Truth Hurdles
Exact match keyphrases can be directly recommended by modeling on click data logs, since they correspond precisely to buyer queries. In contrast, phrase and broad match keyphrases are more difficult to model due to the sparsity and bias of available training signals. It suffers from sparsity as the performance information of phrase/broad match types are only collected when the seller manually chooses them. Although the sellers’ previous picks of successful broad match keyphrases could seem like a good basis for ground truth, they can be heavily affected by the seller set bid amounts and auctioning dynamics [bid_auctioning07], thus making seller-based preference data a bad choice for modeling keyphrase recommendations [dey2025judgejudgeusingllm].
Several biases with click data have been well studied [mishra2025graphexgraphbasedextractionmethod, surveybias, sampleselectionbias, rec4ad, dey2025middlemanbiasadvertisingaligning], these impact any form of supervised training for recommendation. The limited availability of ground truth data leaves a significant portion of the feature space unmarked, requiring the incorporation of counterfactual data so models can learn significant patterns. Moreover, missing label biases [MNAR] affect substantial counterfactual negative signals, as the absence of impressions or clicks doesn’t mean that the match type keyphrases are irrelevant.
@@ I-B2 Active Targeting
Keyphrases of non-exact match type recommended to sellers should be reachable to queries that are searched by buyers. Generated broad/phrase match keyphrases which are open vocabulary cannot reach such queries, though relevant to the item are considered futile leading to wastage of seller bids. [^1] See discussion of Out-of-Vocabulary models [PUSL, simig-etal-2022-open, zest-xml, triangular_bidgen] and targeting in Graphex [mishra2025graphexgraphbasedextractionmethod]. This is exacerbated by the transient queries mentioned in Section I-A1 that require keyphrases to target queries that buyers search for over time. Thus, there should be a way of evaluating the performance of broad matches on transient queries.
@@ I-B4 Execution Requirements
Executing billions of item and query inferences within a few hours is paramount for the timely completion of daily tasks. Given the high memory demands, single-node data aggregation for inference poses a failure risk, while multi-node management elevates engineering expenses. Models need frequent updates for continuous integration of new data points, as discussed in Section I-B2. Deep learning networks like Large Language Models (LLMs) face scrutiny for high execution costs, notably due to their training and inference latency [kaddour2023, mishra2025graphexgraphbasedextractionmethod, 10272546, MLSYS2023_LLMscaling]. Leading LLMs often rely on GPUs, raising acquisition costs and affecting both advertisers’ and platforms’ profit margins.
@@ I-C Scope and Contributions
In this work we only focus on Broad matches instead of Phrase matches [ebayMatchTypes, googleMatchTypes] due to their wide applicability. We use our definition in Section I to streamline the evaluation; the actual implementation of the matching is platform-specific. We aim to tackle the challenges mentioned in Section I-B, which we briefly describe here. We present a novel framework BroadGen for recommending broad match keyphrases to sellers and evaluate its performance. Due to the lack of ground truths per item (see Challenge I-B1), we incorporate relevant buyer queries as input (see Section IV-A) to determine the effective keyphrases for an item. This also helps us target active queries for Challenge I-B2. The Challenge I-B3 implies that a good broad match keyphrase should be neither too generic nor too specific, so that it can achieve a balance between reach and effectiveness. A “too specific” broad match keyphrase will behave like an exact match keyphrase which defeats the purpose. Our approach emphasizes avoiding overly broad suggestions, particularly in Section III-E, with the aim of mitigating inefficiencies within both our systems and the business metrics of sellers. Our framework has been designed to be light and fast with minimal infrastructure costs and is easily deployable in a distributed manner (Challenge I-B4) to cater to millions of sellers with an inventory of over 2.5 billion items daily. We also develop a unique evaluation methodology described in Sections IV-A and  IV-B that enables us to calculate the balance in effectiveness and efficiency of the recommended broad match keyphrases in relation to the search queries it will be matched to. It allows for absence of ground truths and assesses performance over time, emphasizing their adaptability and temporal resilience (Challenge I-B2).
@@ II Related Work
Although broad match keyphrases are a crucial component, the research dedicated to specifically recommending novel broad match keyphrases is sparse. Traditionally, sellers have been responsible for selecting match types, with prior studies [SCHOLZ201996] concentrating on creating keyphrases suitable for exact matches. A simplistic recommendation strategy has involved a rule-based Match Type Curation, which applies heuristics to assess whether an exact match type keyphrase, obtained through the current recall mechanism, has historically succeeded as a broad match keyphrase. This approach, however, faces intrinsic challenges: without historical data indicating that a seller has previously chosen the keyphrase as a broad match type — this method is unable to suggest the keyphrase due to a lack of pertinent performance indicators. Note that in this method the keyphrase is a query in itself.
Figure 3: Example recommendation of keyphrases by the Transfusion using NER tags.
@@
Some studies have focused on optimizing the matching of keyphrases to queries rather than effecting the sellers/advertiser’s selection of keyphrases. The BB-KSM model [LI2022101209] is a stochastic model that utilizes Markov Chain Monte Carlo (MCMC) to initially estimate keyphrase metrics such as impressions and CTR across match types, filling in incomplete data. Subsequently, it designates a match type to each keyphrase based on these metrics with the aim of maximizing a profit objective under budgetary constraints, employing a branch-and-bound algorithm for efficient decision space exploration. Similar optimization strategies have been documented for match types [dar-etal-2009], which recommend using buyer search queries as broad match keyphrases rather than introducing new broad keyphrases. However, with the increasing variety of items and queries in e-commerce, this approach may be inefficient and expensive. The introduction of new queries and changing distributions (refer to Section I-A1) further reduces profitability. Previous studies [radlinski_broder_SIGIR08] apply query substitution methods [rosiejones_querydeletion_sigir03, jones_rey_madani_www06] to connect broad keyphrases with buyer queries by preprocessing frequently searched queries and utilizing predefined substitution mappings.
@@
Transfusion is an eBay’s match type generative model developed based on its in-house NER (Named Entity Recognition) model [xin-etal-2018-learning, palen-michel-etal-2024-queryner] to recommend broad match type keyphrases especially for cold start items. The model generates keyphrases for a specified item utilizing the tokens present in its title. It ensures that these keyphrases align with popular and historically high-performing NER tag combinations within a particular meta category. For instance, as shown in Figure 3, the model identifies that the NER tag combination of product_name and storage_capacity is popular in this category based on the scoring function $f(X)$. The algorithm then construct 3 keyphrases: ‘‘iPhone 5S 16GB’’, ‘‘iPhone 5S 32GB’’, ‘‘iPhone 5S 64GB’’ based on the tokens corresponding to this NER tag combination from the item title. In this work, for comparison, we use a proxy for the rule-based match type curation and the Transfusion model against our model BroadGen.
@@ III-A Terms, Notations and Formulation
Our framework requires a set of existing (buyer search) queries denoted as $Q$ and the item title $t$ for each item to recommend a set of keyphrases denoted as $K$. The goal is for the keyphrase set $K$ to be broad matched to the queries $Q$. Each query $q\in Q$, each keyphrase $k\in K$ and the title $t$ are in the form of a list of tokens. [^2] A proprietary tokenization scheme handles normalization and stop word removal. We formulate the problem as String Clustering on the set of query strings ($Q$), to exploit the commonality within a cluster of queries and generate a representation for each cluster. The underlying intuition is that the cluster representation serves as the broad match keyphrase capable of addressing both the queries within the set and prospective queries beyond it. Our framework has three cores, Input generation (III-B) which generates the input matrix from the set of input queries. The Clustering core (III-C) which groups the queries using the input matrix. The Representation generation (III-D) constructs a representative ordered string for each cluster.
@@ III-B Input Generation
Two input matrices are constructed, Anchor Similarity Matrix $A^{|Q|\times|t|}$ and Query-Query Similarity Matrix $S^{|Q|\times|Q|}$, where $|\cdot|$ determines cardinality. In constructing $A$, a $|t|$-dimensional vector is generated for each query. This vector comprises of binary values, where 1s and 0s indicate the query token’s presence in the title. The construction of $S$ is done by computing the common token count between each pair of queries in $Q$. Each row of $S$ can be considered as a $|Q|$-dimensional vector that represents each query depending on similarity with other queries. Each vector of matrix $S$ is normalized by dividing by its maximum, while each vector of the matrix $A$ is normalized by its sum. The Query-Query Similarity matrix $S$ is transformed into a distance matrix by $D_{ij}=1-S_{ij}$. The core combines both the matrices $A$ and $S$ as shown in
@@
The normalized augmented matrix $\hat{Z}$, provides a vector representation for each query anchoring in relation to the title and acts as a query distance vector when there is insufficient information in the anchor representation.
@@ III-C Clustering Core
The core functionality starts with computing the pairwise distance between each pair of vectors $z_{i},z_{j}\in\hat{Z}$ as demonstrated in
@@ III-D Representation Generation
A representation for each cluster in $C$ depicts a keyphrase that can be extended to queries in the cluster. In the final core, the representation is determined as the set of tokens common among the tokenized queries within each cluster. At times, clusters can be large leading to single or no tokens (in absence of commonality). In these situations, an optimal subset of the cluster is identified whose representation is more than one token. The common tokens from each cluster need to be ordered before concatenating into a string. We employ a graph-based approach to infer an ordering of the common tokens based on their relative positions across the queries. A graph is constructed with all tokens as the vertices; for each co-occurring pair of tokens a directed edge is added indicating local precedence. A topological sort [Tarjan_1976] of the graph using a depth-first search algorithm produces a consistent global ordering of the tokens. This method of determining precedence is rooted in prior work, such as sentence ordering task [prabhumoye-etal-2020-topological]. For each cluster, the common tokens are arranged following the global order and represented as a string.
Figure 4: Example construction of keyphrases by BroadGen from clusters of queries. Illustration shows one cluster per item.
@@ IV-A1 OOT Formulation
To evaluate the effectiveness of broad match keyphrases in addressing the challenge of shifting label distribution, we assess their generalized performance in an Out-of-Time (OOT) [PUSL] setting [OOD2, yu2024surveyevaluationoutofdistributiongeneralization, liu2024robustnessgenerativeinformationretrieval]. Specifically, the algorithms for generating broad match keyphrases use as input the historical queries searched by buyers prior to time $T$, termed as Pre-queries and evaluated against the query distribution observed in the subsequent $T+30$ day period called as Post-queries.